# muography; non-destructive testing; machine learning; image processing; upsampling; semantic segmentation; simulation; cosmic rays

### v1

@@ 3.2 Segmentation Model Results
Using standard evaluation metrics to compare the quality of two images gives a limited interpretation of how the model improves the perception of the features within the muographic images. Additionally, insufficient feature distinction results in the background pixels (attributed to concrete) in the images distorting the evaluation metrics by overwhelming the signal from other classes. In order to evaluate the effect of the upsampling model on each feature type, we trained a second model to perform semantic segmentation of these images, where each pixel in the image is classified as one of the five pre-defined object labels: concrete, rebar grid, tendon duct, air void and unknown. If we were using experimental data we would have to manually label images to retrieve ground truths; however, as we are using Geant4 simulations, the ‘true’ position and composition of each object is known, so we use this to generate our ground truth segmentation images. The segmentation model serves as an independent evaluation metric to assess the performance of the upsampling model. To minimise potential bias and maintain objectivity, this model was trained exclusively on high sampling rate (100-day) muography images, deliberately excluding upsampler outputs from the training process. Model performance is quantified using the Dice coefficient, which measures the overlap between pixel-wise label predictions and ground truth. The five-class segmentation labels were one-hot encoded into discrete binary masks, enabling class-specific overlap calculations. By applying the segmentation model to both original and upsampled muography images, we can assess the upsampler’s effect on each structural class. A comparative analysis of Dice coefficients between original and upsampled images, focusing on the four non-background structural classes, is presented in Figure 3.
The accuracy of the segmentation model for each class is indicated by the performance of the 100-day dataset inputs, showing Dice coefficients for rebar grids, tendon ducts, air voids and ‘unknowns’ of 0.8173, 0.8663, 0.1265 and 0.5132 respectively. In the Dice coefficient difference subplots (in purple) we observe that use of the upsampling model results in an excellent improvement for identification of tendon ducts, a good improvement for rebar and unknowns, and little improvement for air voids. We observe trends similar to that of the SSIM metric in Figure 1, where the upsampling model is most effective at lower
@@
sampling rates before converging with scores similar to the inputs at around 50–70 days of sampling.
We can also perform visual inspection to get a better picture of how the segmentation model is performing. Each of the four panels in Figure 4 show eight versions of a single X-Y plane, seven of different sampling rate versions (a–g) and one as the ground truth (h). The four panels show the original image, outputs of the upsampling model, outputs of the segmentation model and outputs from a concatenation of both models. The upsampled segmentation results (lower right) show a significant improvement at low sampling rates in panels (a) and (b), aligning with trends shown by the Dice coefficients in Figure 3. A key feature of the segmentation model is the lack of vertical rebar present. Looking at the segmentation ground truths in panel (h) we see that the segmentation model is correct in ignoring the vertical rebar indicated by the muography images. This shows that the segmentation model has identified the vertical rebar as a shadowing effect, hence overcoming the z-smearing limitation of muography and ultimately improving the reliability of muography as a technique for accurate concrete analysis.
Figure 3: Average Dice coefficient of the test dataset, calculated at each sampling rate, for each of the four segmentation object labels (rebar grid, tendon duct, void, unknown). Performance of the original input dataset is shown by blue circular points, where upsampled outputs are displayed with red triangular points. The relative improvement of the upsampling model for a particular label is indicated by dice difference below each object plot, shown in purple.
Figure 4: A single X-Y plane image slice for different sampling rates: (a) one day, (b) five days, (c) 10 days, (d) 20 days, (e) 40 days, (f) 60 days, (g) 80 days, (h) ground truth (100-day image for top panels, geometry truth for bottom panels). These eight image versions are displayed as raw input (top left), upsampled (top right), segmented (bottom left), and upsampled and segmented (bottom right). Lilac, blue, red and yellow indicate concrete, rebar, tendon ducts and air voids respectively.
@@ 4 Discussion
The results presented in this study demonstrate promising advancements in applying machine learning techniques to muography image enhancement and analysis. Our findings show that cWGAN-GP models can effectively reduce required sampling times whilst maintaining image quality, with particularly strong performance in feature detection for larger structural elements. The following subsections examine the model performance in detail, addressing both capabilities and limitations for both upsampling and segmentation tasks.
@@ 4.1 Image Upsampling Capabilities
The upsampling model effectively improves image quality as demonstrated by both SSIM and PSNR metrics in Figure 1. Looking at the input images (blue, circular points), both metrics show rapid improvement for sampling rates between one and 35 days. The SSIM follows a logarithmic trend that aligns with the feature detection performance shown in Figure 3 for rebar, tendon ducts and ’unknown’ objects, reflecting its design to quantify perceptual similarity. In contrast, the PSNR, which is calculated as a logarithm of the mean-squared error (MSE), shows a different behaviour: after the initial 30-day period, it increases linearly until approximately 75 days before tending to infinity. This difference arises because PSNR is more sensitive to pixel-level variations, while SSIM captures perceptual changes. The combined analysis suggests that 30–40 days of sampling is sufficient to resolve large-scale features and reduce noise, with additional sampling time primarily improving pixel-level detail that may not significantly impact image perception.
@@
The upsampled outputs (red, triangular points) exhibit systematically higher SSIM and PSNR values compared to the original inputs at lower sampling rates, while following similar overall trends. However, this advantage diminishes as sampling time increases, with upsampled and original images showing identical SSIM performance at around 50 days — the same point at which the corresponding input data (blue, circular points) begins to show significant diminishing returns. This convergence can be understood from the data processing used: each image pixel value corresponds to a specific voxel in its z-plane layer, with longer sampling times allowing more muon hits per voxel. As sampling time increases, these voxel values gradually approach their ‘true’ values, and the 50 day convergence point suggests sufficient sampling has been achieved to closely approximate these true values. The PSNR of the upsampled images converges with its corresponding input images much later, at around 80–85 days, before the input metrics begin to outperform. This is likely caused by the lack of reconstruction of exact pixel-by-pixel noise variations present in the 100-day images, for example shown in Figure 2, where edges and corners exhibit some noise fluctuation that are correspondingly more smooth in the upsampled 1-day image.
@@
Visual inspection of images in Figure 2 and the top panels from Figure 4 can confirm the results exhibited from dataset-wide analysis. At low sampling rates, the noise in the input images is significant enough to drown out most human-perceivable detail, combined with low or non-existent counting statistics around image edges caused by detector acceptance. However, their upsampled counterparts exhibit drastic feature enhancement, combined with a significant reduction in noise to output a smooth but detailed image. While the upsampling of 1-day images demonstrates remarkable improvement, some limitations remain evident — particularly in regions of very low statistics where noise fluctuations dominate. For example, gaps appearing in rebar grids remain unfilled, despite our dataset containing only intact grids, due to the limited context window (4x4 pixels) of the convolutional approach.
@@ 4.2 Feature Segmentation Performance
Figure 3 shows a breakdown of the average segmentation performance for each object feature as a function of sampling rate. The rebar segmentation scores perform well but early convergence occurs between original and upsampled images at around 20 days — notably earlier than the convergence points for SSIM, PSNR, tendon ducts and ‘unknowns’. Also noted is that between 20 and 85 days, the original images slightly outperform the upsampled ones, suggesting the upsampling model is negatively degrading rebar features. Early convergence and outputs being outperformed by inputs suggest the model is washing out some rebar features. This degradation likely stems from the varying dimensions of rebar in our dataset: grids range from 8 to 25 mm, with the thinnest rebar being only four pixels wide at 2 mm pixelisation. The model may be struggling to preserve thin rebar features or mistaking them for shadows of larger ones. The cylindrical grid structure of rebar compounds this challenge, as segmentation in the X-Y plane alone makes edge detection particularly difficult. Visual inspections of samples containing thin rebar (8–10 mm) reveals poor segmentation performance, however a comprehensive breakdown would confirm if this is the true source of the issue.
@@
Tendon ducts perform very well for the input data, however the integration time to achieve the same results with the upsampling model is much improved. This high performance is attributed to the large diameter and pixel volume of the objects. Visual inspection of the dataset also reveals that tendon ducts tend to be placed nearer the edges in low-statistics regions, also attributing to a high relative performance gain — as the large size of the objects mean that they are easily detectable through the high noise. The placement of these ducts near the edges is caused by the concrete sample placement logic (where they are always placed after any rebar grids) which will on average tend to occupy the centres of blocks. Bundling all three tendon duct features together will have also attributed to better object identification.
@@
Air voids are very poorly segmented. The primary causes are their challenging detection in scattering tomographs due to low scattering statistics and their small spherical nature, which constitutes a low pixelwise percentage of the total dataset. Visual inspection of the muography images reveals that most voids greater than 50 mm in diameter can be perceived. For larger voids, at least, the low Dice scores seem to stem from class imbalance. During training, the Dice component of the loss function (Equation 8) is calculated as the mean of the four object components. However, due to the class imbalance — most notably in the case of air voids — the model is not incentivized to improve the correct detection of these pixels. Their low representation in the dataset means they contribute minimally to the loss function, preventing significant improvement in their segmentation. This issue can be addressed by setting custom Dice class weightings such that they reflect the proportion of pixels each class occupies in the test dataset.
@@
The performance of the ‘unknown’ object class is generally satisfactory but leaves room for improvement. This class is particularly challenging to analyse due to its significant variety and density variations. High-density materials, such as uranium and lead tend to exhibit greater smearing, which can negatively impact Dice scores by increasing false positives. Conversely, low-density materials, such as air and water, also contribute to lower Dice scores due to their propensity for false positives. The purpose of this class is to provide the model with a miscellaneous category for objects that do not fall into the predefined categories of rebar, tendon ducts, or air voids. Further analysis of this class, broken down by material composition and size, would provide a clearer understanding of how objects with varying densities perform. Additionally, evaluating models using objects with unseen densities and shapes may provide better insight to the success of this class’ objective.
@@
Visual representations of X-Y plane segmentation outputs are shown in Figure 4, revealing several challenges. The noisy original 1-day image struggles to perform any form of accurate segmentation map — note that it is predicting tendon duct pixels around the edges; the model is ‘guessing’ tendons to exist in low-sample edge regions. Low-sampled upsampled images (panels (a) on the right hand side) which are absent of high noise fluctuations show a much better performance, however seem to struggle to segment all rows of rebar despite being clearly visible. This lack of spatial context limits the model’s ability to accurately interpret the full geometry of the objects, (in complex structures like cylindrical tendon ducts and rebar grids) where straight-edge detection and thicknesses are slightly lacking. In addition to these issues, there are a small number of artefacts exhibited at the top of the images where the shadow of a horizontal tendon duct is present. Many of these are caused by the limited context size of the 4$\times$4 convolutional kernels, which are unable to make predictions based on pixels outwith this area.
4.2.1 A Solution to Smearing Artefacts
@@
We define our inverse imaging problem in muography as the imprecision of a scattering point in a direction orthogonal to the detector planes, caused by the lack of information to place said scattering point on a given muon’s line path. For a detector plane parallel to the ground, the muon path between the two sets of detectors will be along the z-direction, with the increased uncertainty in the z-axis resulting in objects appearing to ‘smear’ along this direction — also referred to as ‘shadows’ or ‘shadowing’. We can see in Figure 4 that the segmented images only contain horizontal rebar, despite the associated muon images appearing to contain a rebar grid of vertical and horizontal rebar. This is due to the segmentation model being trained to reproduce geometric ground truths that were absent of any smearing artefacts, learning the ability to accurately differentiate between shadows and ‘true’ objects, despite a lack of information along the z-axis.
@@
We can investigate this phenomenon further by looking at a concrete block sample side-on, by stacking all segmented X-Y plane images back into a 3-dimensional volume and taking a slice along the X-Z plane. An example of a side-on slice showing the original, upsampled, segmented, and upsampled and segmented versions at different sampling rates are shown in Figure 5. Despite the model not having any context in the z-plane, segmentation outputs (bottom two panels) show remarkable improvements in object clarity when compared to the muographic images. The upsampling model also exhibits a clear improvement to the low sampling rates in panels (a) and (h). However, this ability to reduce smearing effects to accurately distinguish smearing artefacts from actual object geometry would be much improved by using a model input which encompasses all three dimensions.
@@
For simplicity and computational efficiency, our current model processes each X-Y plane layer independently. Whilst effective, this method’s primary limitation is the model’s lack of contextual awareness of a layer’s position within the overall scan and can not utilise information from adjacent layers. However, several methods exist to incorporate 3D context without processing full volumes. These include orthogonal plane processing (where separate models are trained on X-Y, X-Z, and Y-Z planes), multi-layer processing (which analyses small stacks of adjacent layers simultaneously) and patch-based processing (which splits the volume into manageable 3D sub-volumes). These approaches would provide three-dimensional context whilst maintaining reasonable computational requirements and model complexity.

### v2

@@ 3.2 Segmentation Model Results
Using standard evaluation metrics to compare the quality of two images gives a limited interpretation of how the model improves the perception of the features within the muographic images. Additionally, insufficient feature distinction results in the background pixels (attributed to concrete) in the images distorting the evaluation metrics by overwhelming the signal from other classes. In order to evaluate the effect of the upsampling model on each feature type, we trained a second model to perform semantic segmentation of these images, where each pixel in the image is classified as one of the five predefined object labels: concrete, rebar grid, tendon duct, air void, or unknown. If we were using experimental data, we would have to manually label images to retrieve ground truths; however, as we were using Geant4 simulations, the ‘true’ position and composition of each object is known, so we used this to generate our ground truth segmentation images. The segmentation model serves as an independent evaluation metric to assess the performance of the upsampling model. To minimise potential bias and maintain objectivity, this model was trained exclusively on high sampling time (100-day) muography images, deliberately excluding upsampler outputs from the training process. The model performance was quantified using the Dice coefficient, which measures the overlap between pixelwise label predictions and ground truth values. The five-class segmentation labels were one-hot encoded into discrete binary masks, enabling class-specific overlap calculations. By applying the segmentation model to both original and upsampled muography images, we could assess the upsampler’s effect on each structural class. A comparative analysis of Dice coefficients between original and upsampled images, focusing on the four non-background structural classes, is presented in Figure 3.
The accuracy of the segmentation model for each class is indicated by the performance of the 100-day dataset inputs, showing Dice coefficients for rebar grids, tendon ducts, air voids, and ‘unknowns’ of 0.8173, 0.8663, 0.1265, and 0.5132 respectively. For the Dice coefficient difference subplots (in purple), we observe that the use of the upsampling model resulted in an excellent improvement for the identification of tendon ducts, a good improvement for rebar and unknowns, and little improvement for air voids. We observe
@@
Figure 3: Average Dice coefficient of the test dataset, calculated at each equivalent sampling time, for each of the four segmentation object labels (rebar grid, tendon duct, void, unknown). Performance of the original input dataset is shown by blue circular points, where upsampled outputs are displayed with red triangular points. The relative improvement of the upsampling model for a particular label is indicated by Dice difference below each object plot, shown in purple.
trends similar to that of the SSIM metric in Figure 1, where the upsampling model was most effective at lower sampling times before converging with scores similar to the inputs at around 50–70 days of sampling.
We can also perform visual inspection to obtain a better picture of how the segmentation model is performing, here using an image slice which is representative of model performance of on the test dataset. Each of the four panels in Figure 4 show eight versions of a single X–Y plane, seven at different equivalent sampling times (a–g), and one as the ground truth (h). The four panels show the original image, outputs of the upsampling model, outputs of the segmentation model, and outputs from a concatenation of both models. The upsampled segmentation results (lower right) show a significant improvement at low sampling times in panels (a) and (b), aligning with trends shown by the Dice coefficients in Figure 3. A key feature of the segmentation model is the lack of y-aligned rebar features present.
Looking at the segmentation ground truths in panel (h), we see that the segmentation model was correct in ignoring the y-aligned values, as indicated by the muography images. This shows that the segmentation model has identified the y-aligned values as a shadowing effect, hence overcoming the z-smearing limitation of muography and ultimately improving the reliability of muography as a technique for accurate concrete analysis.
Figure 4: A single X–Y plane image slice for different equivalent sampling times: (a) one day, (b) five days, (c) 10 days, (d) 20 days, (e) 40 days, (f) 60 days, (g) 80 days, (h) ground truth (100-day image for top panels, geometry truth for bottom panels). These eight image versions are displayed as raw input (top left), upsampled (top right), segmented (bottom left), and upsampled and segmented (bottom right). Lilac, blue, red, and yellow indicate concrete, rebar, tendon ducts, and air voids, respectively.
@@ 4 Discussion
The results presented in this study demonstrate promising advancements in applying machine learning techniques to muography image enhancement and analysis. Our findings show that cWGAN-GP models can effectively reduce required sampling times while maintaining image quality, with particularly strong performance in feature detection for larger structural elements. The following subsections examine the model performance in detail, addressing both its capabilities and limitations for both upsampling and segmentation tasks.
@@ 4.1 Image Upsampling Capabilities
The upsampling model effectively improved image quality, as demonstrated by both the SSIM and PSNR metrics in Figure 1. Looking at the input images (blue circular points), both metrics showed rapid improvement for equivalent sampling times between one and 35 days. The SSIM followed a logarithmic trend that aligns with the feature detection performance shown in Figure 3 for rebar, tendon ducts, and ’unknown’ objects, reflecting its design to quantify perceptual similarity. In contrast, the PSNR, which is calculated as a logarithm of the mean squared error (MSE), shows a different behaviour: after the initial 30-day period, it increased linearly until approximately 75 days before tending to infinity. This difference arises because the PSNR is more sensitive to pixel-level variations, while the SSIM captures perceptual changes. The combined analysis suggests that 30–40 days of sampling is sufficient to resolve large-scale features and reduce noise, with additional sampling time primarily improving pixel-level detail that may not significantly impact image perception.
@@
The upsampled outputs (red triangular points) exhibited systematically higher SSIM and PSNR values compared to the original inputs at lower sampling times while following similar overall trends. However, this advantage diminished as the sampling time increased, with upsampled and original images showing identical SSIM performance at around 50 days—the same point at which the corresponding input data (blue circular points) began to show significant diminishing returns. This convergence can be understood from the data processing used: each image pixel value corresponds to a specific voxel in its z-plane layer, with longer sampling times allowing more muon hits per voxel. As the sampling time increased, these voxel values gradually approached their ‘true’ values, and the 50-day convergence point suggests sufficient sampling had been achieved to closely approximate these true values. The PSNR of the upsampled images converged with its corresponding input images much later, at around 80–85 days, before the input metrics began to outperform. This was likely caused by the lack of reconstruction of exact pixel-by-pixel noise variations present in the 100-day images, for example, as shown in Figure 2, where the edges and corners exhibited some noise fluctuation that was correspondingly more smooth in the upsampled 1-day image.
@@
Visual inspection of the images in Figure 2 and the top panels from Figure 4 can confirm the results exhibited from dataset-wide analysis. At low sampling times, the noise in the input images was significant enough to drown out most human-perceivable detail, combined with low or non-existent counting statistics around image edges caused by detector acceptance. However, their upsampled counterparts exhibited drastic feature enhancement, combined with a significant reduction in noise to output a smooth but detailed image. While the upsampling of 1-day images demonstrated remarkable improvement, some limitations remain evident—particularly in regions of very low statistics where noise fluctuations dominated. For example, gaps appearing in rebar grids remained unfilled, despite our dataset containing only intact grids, due to the limited context window (4 $\times$ 4 pixels) of the convolutional approach.
@@ 4.2 Feature Segmentation Performance
Figure 3 shows a breakdown of the average segmentation performance for each object feature as a function of equivalent sampling time. The rebar segmentation scores performed well, but early convergence occurred between the original and upsampled images at around 20 days—notably earlier than the convergence points for the SSIM, PSNR, tendon ducts, and ‘unknowns’. Also noted is that between 20 and 85 days, the original images slightly outperformed the upsampled ones, suggesting the upsampling model was negatively degrading rebar features. Early convergence and outputs being outperformed by inputs suggest the model was washing out some rebar features. This degradation likely stems from the varying dimensions of rebar in our dataset: grids range from 8 to 25 mm, with the thinnest rebar being only four pixels wide at 2 mm pixelization. The model may be struggling to preserve thin rebar features or mistaking them for shadows of larger ones. The cylindrical grid structure of the rebar compounds this challenge, as segmentation in the X–Y plane alone makes edge detection particularly difficult. Visual inspections of samples containing thin rebar (8–10 mm) revealed poor segmentation performance; however, a comprehensive breakdown would confirm if this is the true source of the issue.
@@
The tendon ducts performed very well for the input data; additionally, the integration time to achieve the same results with the upsampling model was much improved. This high performance is attributed to the large diameter and pixel volume of the objects. Visual inspection of the dataset also revealed that the tendon ducts tended to be placed nearer the edges in low-statistics regions also attributing to a high relative performance gain—as the large size of the objects means that they are easily detectable through the high noise. The placement of these ducts near the edges is caused by the concrete sample placement logic (where they are always placed after any rebar grids), which will on average tend to occupy the centres of blocks. Bundling all three tendon duct features together will have also resulted in better object identification.
@@
The air voids were very poorly segmented. The primary causes are their challenging detection in scattering tomographs due to low scattering statistics and their small spherical nature, which constitutes a low pixelwise percentage of the total dataset. Visual inspection of the muography images revealed that most voids greater than 50 mm in diameter can be perceived. For larger voids at least, the low Dice scores seem to stem from class imbalance. During training, the Dice component of the loss function (Equation (8)) was calculated as the mean of the four object components. However, due to the class imbalance—most notably in the case of air voids—the model is not incentivised to improve the correct detection of these pixels. Their low representation in the dataset means they contribute minimally to the loss function, preventing significant improvement in their segmentation. This issue can be addressed by setting custom Dice class weightings such that they reflect the proportion of pixels that each class occupies in the test dataset.
@@
The performance of the ‘unknown’ object class was generally satisfactory but leaves room for improvement. This class is particularly challenging to analyse due to its significant variety and density variations. High-density materials, such as uranium and lead, tend to exhibit greater smearing, which can negatively impact Dice scores by increasing false positives. Conversely, low-density materials, such as air and water, also contribute to lower Dice scores due to their propensity for false positives. The purpose of this class is to provide the model with a miscellaneous category for objects that do not fall into the predefined categories of rebar, tendon ducts, or air voids. Further analysis of this class, broken down by material composition and size, would provide a clearer understanding of how objects with varying densities perform. Additionally, evaluating models using objects with unseen densities and shapes may provide better insight to the success of this class’ objective.
@@
Visual representations of the X–Y plane segmentation outputs are shown in Figure 4, revealing several challenges. The noisy original 1-day image struggled to perform any form of accurate segmentation map—note that it is predicting tendon duct pixels around the edges; the model is ‘guessing’ tendons to exist in low-sample edge regions. Low-sampled upsampled images (panels (a) on the right hand side) which were absent of high noise fluctuations showed a much better performance; however, they seemed to struggle to segment all rows of rebar despite being clearly visible. This lack of spatial context limits the model’s ability to accurately interpret the full geometry of the objects (in complex structures like cylindrical tendon ducts and rebar grids), where straight-edge detection and thicknesses are slightly lacking. In addition to these issues, there are a small number of artefacts exhibited at the top of the images where the shadow of an x-aligned tendon duct is present. Many of these are caused by the limited context size of the 4 $\times$ 4 convolutional kernels, which are unable to make predictions based on pixels outwith this area.
A Solution to Smearing Artefacts
@@
We define our inverse imaging problem in muography as the imprecision of a scattering point in a direction orthogonal to the detector planes caused by the lack of information to place said scattering point on a given muon’s line path. For a detector plane parallel to the ground, the muon path between the two sets of detectors will be along the z direction, with the increased uncertainty in the z axis resulting in objects appearing to ‘smear’ along this direction—also referred to as ‘shadows’ or ‘shadowing’. We can see in Figure 4 that the segmented images only contain x-aligned rebar, despite the associated muon images appearing to contain a rebar grid of y- and x-aligned rebar. This is due to the segmentation model being trained to reproduce geometric ground truths that were absent of any smearing artefacts, learning the ability to accurately differentiate between shadows and ‘true’ objects despite a lack of information along the z axis.
@@
We can investigate this phenomenon further by looking at a concrete block sample side-on by stacking all segmented X–Y plane images back into a three-dimensional volume and taking a slice along the X–Z plane. An example of a side-on slice showing the original, upsampled, segmented, and upsampled and segmented versions at different equivalent sampling times is shown in Figure 5. Despite the model not having any context in the z plane, segmentation outputs (bottom two panels) show remarkable improvements in object clarity when compared to the muographic images. The upsampling model also exhibits a clear improvement to the low sampling times in panels (a) and (h). However, this ability to reduce smearing effects to accurately distinguish smearing artefacts from actual object geometry would be much improved by using a model input which encompasses all three dimensions.
@@
For simplicity and computational efficiency, our current model processes each X–Y plane layer independently. While effective, this method’s primary limitation is the model’s lack of contextual awareness of a layer’s position within the overall scan, and it cannot utilise information from adjacent layers. However, several methods exist to incorporate 3D context without processing full volumes. These include orthogonal plane processing (where separate models are trained on X–Y, X–Z, and Y–Z planes), multi-layer processing (which analyses small stacks of adjacent layers simultaneously) and patch-based processing (which splits the volume into manageable 3D sub-volumes). These approaches would provide three-dimensional context while maintaining reasonable computational requirements and model complexity.